# Frequency effects in Linear Discriminative Learning

### v1

@@ 4 Endstate Learning
The first implementation of the DLM in the R (R Core Team,, 2020) package WpmWithLdl (Baayen et al., 2018b, ) estimates the `endstate' of learning. This implementation constructs a mapping between the cue matrix $\mathbf{C}$ and the semantic matrix $\mathbf{S}$ using the pseudo-inverse. Computing the pseudo-inverse as implemented in WpmWithLdl is expensive and prohibitively so for larger datasets. Fortunately, there now exists a very efficient method implemented in the Julia (Bezanson et al.,, 2017) package JudiLing [^4] https://github.com/MegamindHenry/JudiLing.jl that makes use of the Cholesky decomposition (Luo,, 2021; Heitmeier et al.,, 2023). This, together with additional speed-ups due to Julia being in general faster than R and the use of sparse matrices means that JudiLing can handle much larger datasets compared to WpmWithLdl (Luo,, 2021).
@@
Figure 1 illustrates the frequency-free property of EL. It presents the results obtained with an EL model for the dataset introduced in the preceding section, modeling comprehension with the matrix $\mathbf{F}$. The average correlation accuracy for this model is low, at 40.8%; below, we will show how this accuracy can be improved to 83% by increasing the dimensionality of the form vectors.
Figure 1: Endstate learning. The green dots on the horizontal lines at 0 and 1 represent the correlation accuracies@1 for the individual words (counted as correct if the semantic vector most correlated with the predicted semantic vector is the target), and the light blue dots represent the correlation values of words' predicted semantic vectors with their target vectors. The blue line presents the estimated kernel density for log frequency. Target correlation and accuracy are independent of (log) frequency.
To illustrate the absence of a relationship between word frequency and correlation accuracy, we used a binomial generalised linear model [^5] https://github.com/JuliaStats/GLM.jl with a logit link function, modeling the probability of correct recognition as a function of log frequency. Log frequency was not a good predictor of accuracy ($p>0.05$).
In summary, one important advantage of endstate learning is that it is extremely efficient. A second important advantage of this method is that it is frequency-free: it shows what can be learned when frequency of use is not allowed to play a role. In other words, the EL probes the quantitative structure of a dataset, under the assumption that usage can be ignored. Thus, the EL method dovetails well with generative approaches to morphology that deny any role to usage in grammar and work with types only.
@@ 5 Incremental learning
Instead of making use of the efficient method estimating the endstate of learning, we can also learn the mappings incrementally using the Widrow-Hoff learning rule (Widrow and Hoff,, 1960), a form of error-driven learning closely related to the Rescorla-Wagner learning rule (Rescorla and Wagner,, 1972; for a discussion of the Widrow-Hoff learning rule in language research see also Milin et al.,, 2020). Here, the idea is that each time a word is encountered, a learning event occurs, and the mappings between form and meaning are updated in such a way that next time the same word is encountered (if no unlearning occurs intermittently) the mappings will be more accurate. How fast learning takes place is controlled via the learning rate parameter $\eta$. High learning rates lead to fast learning and unlearning, whereas low learning rates result in slower learning.
Incremental learning has the advantage that we can model learning in a frequency-informed way by translating frequencies into learning events. For example, if a word has a frequency of 100, it is presented for learning 100 times. We used the Widrow-Hoff learning rule as implemented in JudiLing to incrementally learn the mapping matrix $\mathbf{F}$. This mapping is now frequency-informed. We experimented with three different learning rates: 0.01, 0.001, and 0.0001. The average correlation accuracy for the three simulations was 15.5%, 14.0% and 10.1% respectively. As can be seen in the left panel of Figure 2, when learning the mappings incrementally according to their frequency, a clear relationship between frequency and accuracy is obtained for all learning rates: the more frequent a word is, the more accurately it is learned. A binomial GLM indicated a highly significant relationship between frequency and accuracy ($p<0.001$ for all learning rates). The lower the learning rate, the steeper the increase in accuracy: accuracy is lower for low-frequency words and higher for high-frequency ones.
@@
Figure 2: Relationship between accuracy and frequency for incremental learning. Left panel: Mapping trained using full frequencies. Predicted accuracy is depicted for three different learning rates ($\eta\in(0.01,0.001,0.0001)$), and the light blue dots present target correlations for $\eta=0.01$. Center panel: Mapping trained using log-transformed frequencies. Right panel: Mapping trained using frequencies scaled by a factor of 100.
@@
There are two disadvantages to this approach, one practical, and the other theoretical. The theoretical problem is that for many datasets, there is no intrinsic order in which words are learned. For the present dataset, which is basically a word frequency list, we do not have available any information about the order in which Dutch speakers encounter these words over their lifetime. It is only for intrinsically ordered data, such as child directed speech in corpora such as CHILDES (MacWhinney,, 2000), that incremental learning comes into its own (see below). The practical problem is that updates with the Widrow-Hoff learning rule are computationally expensive.
For the present dataset, estimating the mapping matrix $\mathbf{F}$ took $\approx$25 minutes on a MacBook Pro (2017) with a 3.1 GHz Quad-Core Intel Core i7 processor, even though the use of bi-grams resulted in a form dimensionality that was already too small to obtain good accuracy. If we would use the better-suited tri-grams (see below for further details), the estimated time for computing the mapping matrix $\mathbf{F}$ increases strongly even in an optimized language such as Julia.
@@
There are two obvious ways in which one might want to avoid the costly computations of WHL. One option is to transform frequencies by taking logarithms (see, e.g., Seidenberg and McClelland,, 1989). The resulting average correlation accuracy is 26.1%. The relationship between frequency and accuracy can be seen in the center panel of Figure 2. While the log transformation does indeed reduce computational costs, it is of course questionable whether such a transformation is justified and realistic. Low frequency words become proportionally more accurate, while high frequency ones become less so. If WHL with empirical frequencies for time-ordered learning events is taken as a gold standard, then a log-transformation distorts our estimates considerably.
@@
A second option is to simply scale down frequencies by dividing them by a fixed number. By applying a ceiling function to the result, we avoid introducing zero frequencies. Training the model using frequencies divided by 100 speeds up the learning to $\approx$12 seconds and does not distort the learning curve (see right panel of Figure 2). The disadvantage of this method is that there are far fewer learning events. As a consequence, words are learned less well. Accordingly, the average correlation accuracy drops to 10.3%.
@@
In summary, it is in principle possible to estimate mapping matrices with incremental learning. This is theoretically highly attractive for data that are intrinsically ordered in learning time (see, e.g., Heitmeier et al.,, 2022, for the modeling of within-experiment learning). For unordered data, some random order can be chosen, but for larger datasets, it would be preferable to have a method that is agnostic about order but nevertheless accounts in a principled way for the consequences of experience for discriminative learning.
6 Non-incremental, yet frequency-informed mappings
@@
We can create two new matrices $\mathbf{C}_{f}$ and $\mathbf{S}_{f}$ where the cue and semantic vectors of the wordforms are repeated according to their frequency count $f_{i}$. We want to find the mapping matrix $\mathbf{F}_{f}$ mapping from $\mathbf{C}_{f}$ to $\mathbf{S}_{f}$. We use a closed-form solution for computing the mapping matrix (see also Baayen et al., 2018a, ):
@@
$=$ $\left(\sum_{i=1}^{m}\sqrt{f_{i}}\mathbf{c}_{i}\left(\sqrt{f_{i}}\mathbf{c}_{i}\right)^{T}\right)^{-1}\left(\sum_{i=0}^{m}\sqrt{f_{i}}\mathbf{c}_{i}\left(\sqrt{f_{i}}\mathbf{s}_{i}\right)^{T}\right)$ (3)
@@
$\mathbf{F}_{f}=\left(\sum_{i=1}^{m}\sqrt{p_{i}}\mathbf{c}_{i}\left(\sqrt{p_{i}}\mathbf{c}_{i}\right)^{T}\right)^{-1}\left(\sum_{i=1}^{m}\sqrt{p_{i}}\mathbf{c}_{i}\left(\sqrt{p_{i}}\mathbf{s}_{i}\right)^{T}\right)$ (5)
Now let $\mathbf{P}\in\mathbb{R}^{m\times m}$ be a diagonal matrix with $p_{ii}=p_{i}$ for $i\in 1,...,m$. Then we can define $\tilde{\mathbf{C}}=\sqrt{\mathbf{P}}\mathbf{C}$ and $\tilde{\mathbf{S}}=\sqrt{\mathbf{P}}\mathbf{S}$ so that $\tilde{\mathbf{c}}_{i}=\sqrt{p_{i}}\mathbf{c}_{i}$ and $\tilde{s}_{i}=\sqrt{p_{i}}\mathbf{s}_{i}$. Then we have
@@
Practically, this means that we can first weigh $\mathbf{C}$ and $\mathbf{S}$ with the pertinent frequencies to obtain $\tilde{\mathbf{C}}$ and $\tilde{\mathbf{S}}$. We can then use the closed-form solution (making use of algorithms such as Cholesky decomposition) to obtain frequency-informed mappings between these two matrices. [^6] Note that while we here work with a learning rule for continuous vectors, FIL is also applicable to mappings between discrete vectors, as in Rescorla-Wagner learning (Rescorla,, 1967) which is used in Naive Discriminative Learning (NDL, Baayen et al.,, 2011): instead of having to impose some random order on the learning events, the expected value irrespective of order can be estimated using FIL.
@@
6.1 Low-dimensional modeling
We first consider modeling studies using the low-dimensional vectors that we used in the preceding sections. The low dimensionality of these vectors is motivated by the wish to avoid long computation times for WHL for these exploratory studies.
Figure 3 shows the relationship between log-transformed frequency and accuracy predicted by a logistic GLM regressing the correctness of FIL responses on log frequency. Accuracy@k is set to 1 if a word's predicted semantic vector is among the $k$ predicted vectors that are most correlated with the targeted semantic vector, and to 0 otherwise.
Figure 3: Frequency-informed learning. The red line presents the model's predictions when a success is defined as the predicted vector being the closest to its gold standard target vector in terms of correlation (accuracy@1). The purple line represents model predictions when a success is defined as the correlation being among the top 10 (accuracy@10). The blue line visualizes the estimated density of the log-transformed frequencies. The green dots represent the successes and failures for accuracy@1. The light blue dots represent for each word the correlation of the predicted and gold-standard semantic vectors.
Correlation and accuracy@1 increase for higher frequency, as required. When comparing accuracies with the frequency distribution depicted in the same plot, we can also see that there is a large number of low frequency words with very low (predicted) accuracy. Although for most words, the correlations are relatively high, the overall accuracy@1 is low, at 9.9%. When we relax our criterion for accuracy, use accuracy@10, counting a predicted vector as correct if the target vector is among the 10 closest semantic vectors, we see that the accuracy starts to rise earlier, but there is still a significant portion of words for which even accuracy@10 is zero.
The relatively large number of words with very low accuracy raises the question of whether accuracy can be improved by using log-transformed frequencies for FIL. Figure 4 clarifies that accuracies increase for lower-frequency words, but decrease somewhat for higher-frequency words. The average accuracy@1 is accordingly higher at 34.3%.
@@ 6.1 Low-dimensional modeling
The upper half of Table 1 provides an overview of the accuracies@1 for different combinations of learning (incremental/frequency-informed) and kind of frequency used (untransformed, scaled, or log-transformed). Here, we observe first of all that endstate learning offers the highest accuracy (40.8%), followed by log-frequency informed learning (34.3%).
low-dimension simulations
@@
Incremental learning log-frequencies 26.1% 57.6%
@@
Log-frequency-informed learning 34.3% 71.2%
high-dimension simulations
@@
Table 1: Comparison of average and frequency-weighted accuracy@1 (the term ``frequency-weighted accuracy'' is introduced in Section 6.4) across simulation studies. The small dataset contained 2,638 word forms, the large dataset 13,669.
@@
Figure 5: Comparison of methods. GLM-predicted Accuracy@1 with frequency-informed learning is plotted as a black solid line: The left panel compares methods based on log-frequencies, the center panel compares methods based on scaled frequencies and the right panel compares incremental learning with different learning rates. Incremental learning with scaled frequencies or with a very low learning rate ($\eta=0.0001$) is closest to frequency-informed learning.
@@
6.2 High-dimensional modeling
Importantly, the mappings that we used thus far are suboptimal: the dimensionality of the semantic vectors was small and the use of bi-grams for the form vectors often shows underwhelming performance (see Heitmeier et al.,, 2021, for further discussion of the underlying reasons). While opting for low dimensionality decreased the computational costs for incremental learning immensely and was therefore necessary for comparing methods, we now proceed to investigate the accuracy of frequency-informed learning for larger, more discriminative cue matrices. To this end, we employed letter trigrams resulting in a form vector dimensionality of 1776 (compared to 360 previously).

### v2

@@ 3.2 Endstate Learning
Computing the pseudo-inverse as implemented in WpmWithLdl is expensive and prohibitively so for larger datasets. Fortunately, there now exists a very efficient method implemented in the Julia (Bezanson et al.,, 2017) package JudiLing [^4] https://github.com/MegamindHenry/JudiLing.jl that makes use of the Cholesky decomposition (Luo,, 2021; Heitmeier et al., 2023b, ). This, together with additional speed-ups due to Julia being in general faster than R and the use of sparse matrices means that JudiLing can handle much larger datasets compared to WpmWithLdl (Luo,, 2021).
@@
Figure 1: Endstate learning. The green filled dots on the horizontal lines at 0 and 1 represent the correlation accuracies@1 for the individual words (counted as correct if the semantic vector most correlated with the predicted semantic vector is the target), and the light pink circles represent the correlation values of words' predicted semantic vectors with their target vectors. The dark blue dotted line presents the estimated kernel density for log frequency.
There is no discernible relationship between Log Frequency and correlation/accuracy for endstate learning.
Figure 1 illustrates the frequency-free property of EL. It presents the results obtained with an EL model with low-dimensional form vectors for the small dataset introduced in the preceding section, modeling comprehension with the matrix $\mathbf{F}$. The average correlation accuracy for this model is low, at 40.8%; below, we will show how this accuracy can be improved to 83% by increasing the dimensionality of the form vectors.
To illustrate the absence of a relationship between word frequency and correlation accuracy, we used a binomial generalised linear model [^5] https://github.com/JuliaStats/GLM.jl with a logit link function, modeling the probability of correct recognition as a function of log frequency (here and in all later analyses, before log-transformation, a backoff value of 1 was added to word frequencies). Log frequency was not a good predictor of accuracy ($p=0.9772$).
In summary, one important advantage of endstate learning is that it is can be computed very quickly: on a MacBook Pro (2017) with a 3.1 GHz Quad-Core Intel Core i7 processor and for the small dataset with low-dimensional form vectors it takes about 20 milliseconds. A second important advantage of this method is that it is frequency-free: it shows what can be learned when frequency of use is not allowed to play a role. In other words, the EL probes the quantitative structure of a dataset, under the assumption that usage can be ignored. Thus, the EL method dovetails well with generative approaches to morphology that deny any role to usage in grammar and work with types only.
@@ 3.3 Incremental learning
Instead of making use of the efficient method estimating the endstate of learning, we can also learn the mappings incrementally using the Widrow-Hoff learning rule (Widrow and Hoff,, 1960), a form of error-driven learning closely related to the Rescorla-Wagner learning rule (Rescorla and Wagner,, 1972; for a discussion of the Widrow-Hoff learning rule in language research see also Milin et al.,, 2020). Here, the idea is that each time a word is encountered, a learning event occurs, and the mappings between form and meaning are updated in such a way that next time the same word is encountered (if no unlearning occurs intermittently) the mappings will be more accurate. To be precise, instead of obtaining the mapping matrix $\mathbf{F}$ via equation 1, it is learned gradually via the following equation:
$\mathbf{F}_{t+1}=\mathbf{F}_{t}+\mathbf{c}_{t}^{T}\cdot(\mathbf{s}_{t}-\hat{\mathbf{s}}_{t})\cdot\eta$
where $\mathbf{F}_{t}$ is the state of the matrix $\mathbf{F}$ at learning step $t$, $\mathbf{c}_{t}$ and $\mathbf{s}_{t}$ are the form and semantic vectors of the wordform encountered at $t$, and $\hat{\mathbf{s}}_{t}=\mathbf{c}_{t}\cdot\mathbf{F}_{t}$. How fast learning takes place is controlled via the learning rate parameter $\eta$. High learning rates lead to fast learning and unlearning, whereas low learning rates result in slower learning and less unlearning.
@@
Figure 2: Relationship between accuracy and frequency for incremental learning. Left panel: Mapping trained using full frequencies. Predicted accuracy is depicted for three different learning rates ($\eta\in\{0.01,0.001,0.0001\}$), and the light pink circles present target correlations for $\eta=0.01$. Center panel: Mapping trained using log-transformed frequencies. Right panel: Mapping trained using frequencies divided by a factor of 100.
While there is a strong relationship between log frequency and accuracy/correlation when training on full frequencies and scaled frequencies, this relationship is attenuated when training on log-transformed frequencies.
Incremental learning has the advantage that we can model learning in a frequency-informed way by translating frequencies into learning events. For example, if a word has a frequency of 100, it is presented for learning 100 times. We used the Widrow-Hoff learning rule as implemented in JudiLing to incrementally learn the mapping matrix $\mathbf{F}$. This mapping is now frequency-informed. We experimented with low-dimensional form vectors and three different learning rates: 0.01, 0.001, and 0.0001, modelling words in the small dataset. The average correlation accuracy for the three simulations was 15.5%, 14.0% and 10.1% respectively. As can be seen in the left panel of Figure 2, when learning the mappings incrementally according to their frequency, a clear relationship between frequency and accuracy is obtained for all learning rates: the more frequent a word is, the more accurately it is learned. A binomial GLM indicated a highly significant relationship between frequency and accuracy ($p<0.001$ for all learning rates). The lower the learning rate, the steeper the increase in accuracy: accuracy is lower for low-frequency words and higher for high-frequency ones.
@@
There are two disadvantages to this approach, one practical, and the other theoretical. The theoretical problem is that for many datasets, there is no intrinsic order in which words are learned. For the present dataset, which is basically a word frequency list, we do not have available any information about the order in which Dutch speakers encounter these words over their lifetime. It is only for intrinsically ordered data, such as child directed speech in corpora such as CHILDES (MacWhinney,, 2014), that incremental learning comes into its own (see Section 5). The practical problem is that updates with the Widrow-Hoff learning rule are computationally expensive. For the present dataset, estimating the mapping matrix $\mathbf{F}$ took $\approx$25 minutes on a MacBook Pro (2017) with a 3.1 GHz Quad-Core Intel Core i7 processor, even though the use of bi-grams resulted in a form dimensionality that was already too small to obtain good accuracy. If we would use the better-suited tri-grams (i.e. the high-dimensional form vectors described in Section 3.1.2), the estimated time for computing the mapping matrix $\mathbf{F}$ increases strongly even in an optimized language such as Julia.
@@
The computational cost of WHL may be alleviated in (at least) two ways. One option is to transform frequencies by taking logarithms (see, e.g., Seidenberg and McClelland,, 1989). The resulting average correlation accuracy is 26.8%. The relationship between frequency and accuracy can be seen in the center panel of Figure 2. While the log transformation does indeed reduce computational costs, it is questionable whether such a transformation is justified and realistic. Low frequency words become proportionally more accurate, while high frequency ones become less so. If WHL with empirical frequencies for time-ordered learning events is taken as a gold standard, then a log-transformation distorts our estimates considerably.
@@
A second option is to simply scale down frequencies by dividing them by a fixed number. By applying a ceiling function to the result, we avoid introducing zero frequencies. Training the model using frequencies divided by 100 speeds up the learning to $\approx$12s and does not distort the learning curve (see right panel of Figure 2). The disadvantage of this method is that there are far fewer learning events. As a consequence, words are learned less well. Accordingly, the average correlation accuracy drops to 10.3%.
@@
In summary, it is in principle possible to estimate mapping matrices with incremental learning. This is theoretically highly attractive for data that are intrinsically ordered in learning time (see, e.g., Heitmeier et al., 2023a, , for the modeling of within-experiment learning). For unordered data, some random order can be chosen, but for larger datasets, it would be preferable to have a method that is agnostic about order but nevertheless accounts in a principled way for the consequences of experience for discriminative learning.
3.4 Non-incremental, yet frequency-informed mappings
@@
We can create two new matrices $\mathbf{C}_{f}$ and $\mathbf{S}_{f}$ where the cue and semantic vectors of the wordforms are repeated according to their frequency count $f_{i}$. We want to find the mapping matrix $\mathbf{F}_{f}$ mapping from $\mathbf{C}_{f}$ to $\mathbf{S}_{f}$. We use the following solution for computing the mapping matrix (see also Baayen et al., 2018a, ):
@@
$=$ $\left(\sum_{i=1}^{m}\sqrt{f_{i}}\mathbf{c}_{i}\left(\sqrt{f_{i}}\mathbf{c}_{i}\right)^{T}\right)^{-1}\left(\sum_{i=1}^{m}\sqrt{f_{i}}\mathbf{c}_{i}\left(\sqrt{f_{i}}\mathbf{s}_{i}\right)^{T}\right)$ (4)
@@
$\mathbf{F}_{f}=\left(\sum_{i=1}^{m}\sqrt{p_{i}}\mathbf{c}_{i}\left(\sqrt{p_{i}}\mathbf{c}_{i}\right)^{T}\right)^{-1}\left(\sum_{i=1}^{m}\sqrt{p_{i}}\mathbf{c}_{i}\left(\sqrt{p_{i}}\mathbf{s}_{i}\right)^{T}\right).$ (6)
Now let $\mathbf{P}\in\mathbb{R}^{m\times m}$ be a diagonal matrix with $p_{ii}=p_{i}$ for $i\in 1,...,m$. Then we can define $\tilde{\mathbf{C}}=\sqrt{\mathbf{P}}\mathbf{C}$ and $\tilde{\mathbf{S}}=\sqrt{\mathbf{P}}\mathbf{S}$ so that $\tilde{\mathbf{c}}_{i}=\sqrt{p_{i}}\mathbf{c}_{i}$ and $\tilde{\mathbf{s}}_{i}=\sqrt{p_{i}}\mathbf{s}_{i}$. Then we have
@@
Practically, this means that we can first weigh $\mathbf{C}$ and $\mathbf{S}$ with the pertinent frequencies to obtain $\tilde{\mathbf{C}}$ and $\tilde{\mathbf{S}}$. We can then use the solution in equation 1 (making use of algorithms such as Cholesky decomposition) to obtain frequency-informed mappings between these two matrices. [^6] Note that while we here work with a learning rule for continuous vectors, FIL is also applicable to mappings between discrete vectors, as in Rescorla-Wagner learning (Rescorla,, 1967) which is used in Naive Discriminative Learning (NDL, Baayen et al.,, 2011): instead of having to impose some random order on the learning events, the expected value irrespective of order can be estimated using FIL.
@@
3.4.1 Low-dimensional modeling
We first consider modeling studies using the low-dimensional vectors that we used in the preceding sections together with the small dataset. We chose this low dimensionality in order to avoid long computation times for WHL for these exploratory studies.
Figure 3: Frequency-informed learning. The red solid line presents the predictions of a GLM when a success is defined as the predicted vector being the closest to its gold standard target vector in terms of correlation (accuracy@1). The light blue dashed line represents model predictions when a success is defined as the correlation being among the top 10 (accuracy@10). The dark blue dotted line visualizes the estimated density of the log-transformed frequencies. The green filled dots represent the successes and failures for accuracy@1. The light pink circles represent for each word the correlation of the predicted and gold-standard semantic vectors.
There is a strong relationship between log frequency and correlation/accuracy, and the GLM-predicted accuracy@10 is shifted to the left, i.e. accuracy@10 rises for lower frequencies.
Figure 3 shows the relationship between log-transformed frequency and accuracy predicted by a logistic GLM regressing the correctness of FIL responses on log frequency. Accuracy@k is set to 1 if a word's target semantic vector is among the $k$ target vectors that are most correlated with the predicted semantic vector, and to 0 otherwise.
Correlation and accuracy@1 increase for higher frequency, as required. When comparing accuracies with the frequency distribution depicted in the same plot, we can also see that there is a large number of low frequency words with very low (predicted) accuracy. Although for most words, the correlations are relatively high, the overall accuracy@1 is low, at 9.9%. When we relax our criterion for accuracy, using accuracy@10, counting a predicted vector as correct if the target vector is among the 10 closest semantic vectors, we see that the accuracy starts to rise earlier, but there is still a significant portion of words for which even accuracy@10 is zero.
@@ 3.4.1 Low-dimensional modeling
When FIL is trained with log-transformed frequencies, lower-frequency words are recognized more accurately, but higher-frequency words less accurately.
The relatively large number of words with very low accuracy raises the question of whether accuracy can be improved by using log-transformed frequencies for FIL. Figure 4 clarifies that accuracies increase for lower-frequency words, but decrease somewhat for higher-frequency words. The average accuracy@1 is accordingly higher at 34.6%.
The upper half of Table 1 provides an overview of the accuracies@1 for different combinations of learning (incremental/frequency-informed) and kind of frequency used (untransformed, scaled, or log-transformed). Here, we observe first of all that endstate learning offers the highest accuracy (40.8%), followed by log-frequency informed learning (34.6%).
low-dimensional form vectors
@@
Incremental learning log-frequencies 26.8% 65.4%
@@
Log-frequency-informed learning 34.6% 71.3%
high-dimensional form vectors
@@
Table 1: Comparison of average and frequency-weighted accuracy@1 (the term ``frequency-weighted accuracy'' is introduced in Section 3.4.4) across simulation studies. The small dataset contained 2,638 word forms, the large dataset 13,669. When working with low-dimensional form vectors, the small dataset was used throughout.
Figure 5: Comparison of methods. GLM-predicted Accuracy@1 with frequency-informed learning is plotted as a black line: The left panel compares methods based on log-frequencies, the center panel compares methods based on scaled frequencies and the right panel compares incremental learning with different learning rates. Incremental learning with scaled frequencies or with a very low learning rate ($\eta=0.0001$) is closest to frequency-informed learning.
@@
3.4.2 High-dimensional modeling
Importantly, the mappings that we used thus far are suboptimal: the dimensionality of the semantic vectors was small and the use of bi-grams for the form vectors often shows underwhelming performance (see Heitmeier et al.,, 2021, for further discussion of the underlying reasons). While opting for low dimensionality decreased the computational costs for incremental learning immensely and was therefore necessary for comparing methods, we now proceed to investigate the accuracy of frequency-informed learning for larger, more discriminative cue matrices. To this end, we next experimented with the high-dimensional form vectors, still making use of the small dataset.